OpenStudy (anonymous):

For the function f(x) = 2 − 3x, find f(−4). −10 −5 12 14

OpenStudy (anonymous):

Substitute -4 for x and evaluate.\[f \left( -4 \right)=2-3\left( -4 \right) = \text{?}\]

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

No. What is -3 x -4?

OpenStudy (anonymous):

12 Woops.

OpenStudy (anonymous):

Great. Now add 2. What do you get?

OpenStudy (anonymous):

12+2=14 JKJK 21

OpenStudy (anonymous):

14 is correct. Good job!